# Quantifying the Impact of Incompleteness on Identifying and Interpreting Galaxy Protocluster Populations with the TNG-Cluster Simulation

### v1

@@ 3.3 Evaluating Completeness Limits for Recovering the True Highest Galaxy Density Peak
In Fig. 8 we explore the stellar mass and SFR completeness limits required to recover the location of the baseline highest galaxy density peak. The left column presents the 3D separation between the baseline and stellar mass-limited highest galaxy density peaks, with the top, middle, and bottom rows showing the data binned by cluster halo mass at $z=0$. Separations are plotted as a function of redshift, with three stellar mass thresholds, ranging from $10^{9.0}$ to $10^{10.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. While the scatter is significant, the separations generally correlate with the cluster halo mass at $z=0$, with the median separation increasing for more massive $M_{\rm{200}}^{z=0}$ bins. We also find that separations are sensitive to changes in the minimum stellar mass, as even a $0.5$ dex increase from the baseline stellar mass (i.e., ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) results in separations $\gtrsim 1$ pMpc at $z\geq 3$.
@@
In the right column of Fig. 8, we show the 3D separation between the baseline and SFR-limited highest galaxy density peaks, with the SFR-limited sample restricted to ${M}_{\star}>10^{9.0}\leavevmode\nobreak\ {\rm M}_{\odot}$ and ranging from $0.1$ to $5.0\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$. In general, we find that the correlation between the cluster halo mass at $z=0$ and the magnitude of the median separations is relatively weak. Additionally, while the scatter is large, the median separations generally exceed $0.5$ pMpc at $z>2$ and increase with more restrictive SFR thresholds.
@@
In Fig. 9, we plot the baseline highest density peak recovery fraction as a function of highest density peak-centric separation. The left column shows the cumulative recovery fraction, binned by redshift from $z=1$ to $z=5$, for the separation between the baseline highest galaxy density peak and ${M}_{\star}$-limited highest galaxy density peaks for stellar mass thresholds ranging from $10^{9.0}$ to $10^{10.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. For ${M}_{\star}>10^{9.0}\leavevmode\nobreak\ {M}_{\star}$, approximately $40\%$ of the sample recovers the baseline density peak for separations less than $1$ pMpc; however, this drops to $20\%$ for peaks traced by protocluster populations with ${M}_{\star}>10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$.
@@
The right column of Fig. 9 shows the cumulative recovery fraction for the separation between the baseline highest galaxy density peak and the highest galaxy density peaks traced by SFR-limited subpopulations with SFR thresholds ranging from $0.1$ to $5.0\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$. For all of the SFR thresholds explored, the recovery rate is consistently around $40\%$ at $z>2$ for scales less than 1 pMpc. Given that this sample is constrained to ${M}_{\star}>10^{9.0}\leavevmode\nobreak\ {\rm M}_{\odot}$, it suggests that stellar mass, rather than SFR, is the primary driver of the discrepancy between the locations of the observationally-limited and baseline highest galaxy density peaks. Overall, these results show that the location of the highest concentration of galaxies for protocluster populations complete down to ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ is generally not recovered on scales less than 1 pMpc at $z>2$.
@@
Figure 10: Ratio of the average SFR (top row) and stellar mass (bottom row) within $R_{10}$ centered on the baseline highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$), compared to the average SFR and stellar mass in the surrounding volume outside of $R_{10}$. The columns show these ratios binned by cluster mass at $z=0$. A box-and-whisker plot visualizes the distribution, with vertical boxes representing the interquartile range (25th to 75th percentiles), horizontal bars and circles indicating the median and mean ratios, and whiskers extending to 1.5 times the interquartile range, covering approximately $95\%$ of the distribution. While the mean ratios are generally above unity, they are sensitive to outliers. The median ratios, which are robust to outliers, show that the average SFR and stellar mass near the highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$) are lower than in the protocluster as a whole. This suggests that regions with the highest galaxy densities are not necessarily associated with enhanced star formation or stellar mass growth.
Figure 11: Projected density distribution for the progenitor of the most massive $z=0$ galaxy cluster as a function of redshift. The top row shows the galaxy density map, while the middle and bottom rows display the stellar mass and dark matter weighted density maps from TNG-Cluster, respectively. The location of the highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$) is marked by a black cross, while the highest stellar mass ($\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$) and dark matter ($\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$) peaks are indicated by black plus and hexagon symbols, respectively. In this cluster the locations of these highest density peaks overlap at $z\leq 2$ but are separated by $1-3$ pMpc at earlier times. This suggests that the highest galaxy density peak does not necessarily coincide with the highest concentrations of stellar mass and dark matter in protoclusters.
@@
To examine the impact of potentially missing the highest galaxy density peak in observations, we explore the properties of galaxies (e.g., SFRs or stellar masses) inside and outside this region in our simulated protocluster population. We compare these properties by calculating the ratio of the average SFR and stellar mass inside and outside $R_{10}$, measured relative to the location of the highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$). Ratios close to unity imply that the galaxies in the highest density peak have properties on average similar to those of galaxies in the rest of the protocluster. Ratios below or above unity indicate that the galaxies within the highest density peak have properties that are either depressed or elevated relative to the remainder of the protocluster.
@@
In Fig. 10 we show the ratio of the average SFR and stellar mass as a function of redshift for galaxies within $R_{10}$ centered on the highest galaxy density peak, compared to those beyond this region. The data is binned by cluster halo mass at $z=0$ and presented as a box-and-whisker plot, with the vertical boxes representing the interquartile range (25th to 75th percentiles), and the horizontal bars and circles showing the median and mean of the distribution. The whiskers extend to 1.5 times the interquartile range, encompassing approximately $95\%$ of the distribution.
The top row of Fig. 10 shows that for the average SFR within the highest galaxy density peak the mean of this distribution is consistently above unity, regardless of the cluster mass at $z=0$, compared to the remainder of the protocluster. As this trend can be influenced by outliers, we also examine the median of the distribution. We find that, regardless of the cluster mass at $z=0$, the median SFR within the highest galaxy density peak is below unity for $z>2$, indicating that the average SFR in this region is depressed relative to the rest of the protocluster.
In the bottom row of Fig. 10 we show the ratio of the average stellar mass within the highest density peak compared to the average stellar mass of the protocluster. The mean of this distribution is consistently above unity for all redshifts, with very little dependence on the halo mass of the cluster at $z=0$. Meanwhile, we find that the median stellar mass ratio is generally below unity at $z>3$ but increases with redshift toward $z=0$. Additionally, we observe a mild dependence on the host halo mass of the cluster at $z=0$, with progenitors of the most massive clusters reaching unity earlier. These results imply that the regions with the highest concentration of galaxies in protocluster populations complete down to ${M}_{\star}\sim 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ are typically not regions of accelerated galaxy evolution relative to the remainder of the protocluster.
@@
Figure 12: Cumulative recovery fraction of the baseline stellar mass and dark matter highest density peaks, binned by redshift and measured relative to the highest galaxy density peaks traced by various ${M}_{\star}$- and SFR-limited subpopulations. The first two rows show separations from the baseline stellar mass (first row) and dark matter (second row) highest density peaks to the highest galaxy density peak traced by ${M}_{\star}$-limited subpopulations. The third and fourth rows show the same for SFR-limited subpopulations. Columns represent increasingly restrictive stellar mass ($10^{8.5}$ to $10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$) and SFR (0.10.10.10.$10.0\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$) thresholds. Regardless of the selected ${M}_{\star}$ or SFR threshold, the highest stellar mass and dark matter density peaks are recovered in $\lesssim 40\%$ of cases with accuracies within 1 pMpc ($2.1-2.6\arcmin$) for $2<z\leq 5$. However, in the vast majority of cases, these peaks are only recovered at larger separations of 14141–41 – 4 pMpc.
@@
The results presented in the previous section appear to conflict with findings from observed protocluster studies, which suggest that the densest regions of galaxy protoclusters are sites of accelerated galaxy growth and prodigious star formation (e.g., Steidel et al., 2005; Dannerbauer et al., 2014; Hayashi et al., 2016; Miller et al., 2018; Shimakawa et al., 2018a). However, this discrepancy may stem from the specific region of the galaxy protocluster being probed. Thus far we have considered only the region of the highest galaxy density peak, which is commonly used to characterize the density field of observed protocluster galaxies. Alternatively, the highest matter density peak could be examined by weighting the galaxy density field according to the stellar and halo masses of the protocluster galaxies. To achieve a balanced weighting scheme that prevents either high-mass or low-mass galaxies from dominating the density field, we define the weights applied to the 3D histogram as $w=M^{\alpha}+C$, where $M$ represents either the stellar mass or dark matter component of each galaxy. We set $\alpha=0.5$ to ensure a balanced contribution from low-mass and high-mass galaxies, while the constant $C=0.5$ establishes a baseline weight [^2] In practice, $\alpha$ only affects the map in the range between 0 and 1, with $\alpha\leq 0$ returning the original galaxy density map, and $\alpha\geq 1$ excluding contributions from low-mass galaxies, making the map solely reflect the location of the most massive galaxy..
@@
In Fig. 11 we present the projected galaxy density (top row), stellar mass density (middle row), and halo mass density (bottom row) for the progenitor of the most massive galaxy cluster at $z=0$. The columns display the projected density maps at five distinct redshifts, ranging from $z=1$ to $z=5$. We overlay the location of the highest density peak for each density field, marking the highest galaxy density peak with a black cross, the highest stellar mass peak with a black plus sign, and the highest dark matter mass density peak with a black hexagon. For this specific test case we find that above $z>2$ the location of the highest galaxy density peak does not always coincide with the location of the highest stellar mass or halo mass density peaks. Observationally, the opposite is typically observed, for instance, Koyama et al. (2021) and Polletta et al. (2021) find that the densest region of the Planck-selected protocluster PHz G237.01+42.50 contains the most massive members.
@@
Figure 13: Box-and-whisker plot of the ratio of the average SFR (top row) and stellar mass (bottom row) within $R_{10}$ centered on the highest density peak for dark matter (blue boxes, $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$), stellar mass (orange boxes, $\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$), and galaxies (yellow boxes, $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$), compared to the average SFR and stellar mass in the surrounding volume outside $R_{10}$. The columns display these ratios in bins of cluster mass at $z=0$. In contrast to the ratios for $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$, the matter density peaks exhibit ratios significantly above unity for both SFR and stellar mass, suggesting that galaxies in these regions experience accelerated growth and enhanced star formation compared to the protocluster as a whole.
@@
The first two rows of Fig. 12 show the cumulative matter density recovery fractions measured from the baseline stellar mass (first row) and dark matter (second row) highest density peaks to the highest galaxy density peak traced by four ${M}_{\star}$-limited subpopulations. Moving from left to right, the stellar mass thresholds increase from $10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ (i.e., the baseline stellar mass established in this study) to $10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$. Regardless of the stellar mass threshold, the true highest stellar mass and dark matter density peaks are only recovered in $\lesssim 40\%$ of cases within an accuracy of 1.0 pMpc ($\sim 2.1-2.6\arcmin$) at $2<z\leq 5$. However, at $z=5$, the recovery fraction of the highest stellar mass density peak increases with more restrictive stellar mass thresholds when considering separations within 0.5 pMpc ($1.3\arcmin$). Specifically, the recovery fraction rises from $15\%$ for ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ to $30\%$ for ${M}_{\star}>10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$, suggesting that samples containing more massive galaxies better trace the highest stellar mass density peak on small scales.
A similar trend, but in the opposite direction, is observed for the recovery fraction of dark matter density peak at $z=5$ when considering separations with accuracies within 0.5 pMpc ($1.3\arcmin$). In this case, the recovery fraction decreases from $30\%$ for ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ to $25\%$ for ${M}_{\star}>10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$. Nevertheless, in most cases, the matter density peaks are only recovered at separations between 1 and 4 pMpc, suggesting that galaxy density peaks inferred from ${M}_{\star}$-limited populations do not generally trace the true matter density peaks.
@@
The last two rows of Fig. 12 show the cumulative matter density recovery fractions measured from the baseline stellar mass (third row) and dark matter (fourth row) highest density peaks to the highest galaxy density peak traced by four SFR-limited subpopulations. Moving from left to right, the SFR thresholds increase from $0.1$ to $10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$. The SFR-limited results, which are not additionally constrained by stellar mass as in Fig. 9, are quite similar to the stellar mass-limited results. Specifically, matter density peaks are recovered in only $\sim 40\%$ of cases within an accuracy of 1.0 pMpc ($\sim 2.1-2.6\arcmin$) for $2<z\leq 5$, while the majority of samples recover the highest matter density peaks only at separations between 1 and 4 pMpc for $z>2$.
The cumulative recovery fraction results from Fig.12 suggest that regardless of the stellar mass or SFR limitations of the protocluster galaxy population, the true matter density peaks are only recovered in $\sim 40\%$ of cases with an accuracy of 1.0 pMpc ($\sim 2.1-2.6\arcmin$) for $2<z\leq 5$. However, these results depend on the weighting scheme used in the matter density maps. The chosen value of $\alpha=0.5$ provides a balanced approach, with $\alpha\leq 0$ reproducing the baseline galaxy density field and $\alpha\geq 1$ being overly influenced by the location of the most massive galaxy. Slight deviations from this parameter choice would only marginally affect the agreement between the galaxy and matter density peaks. Despite this caveat, these results suggest that the highest density peaks traced by galaxies do not always coincide with the location of matter density peaks, likely due to the fact that the baseline highest galaxy density peak does not always coincide with the most massive galaxies or BCG progenitors, which are generally more reliable tracers of the matter density field (see Fig. 6).

### v2

@@ 3.3 Evaluating Completeness Limits for Recovering the True Highest Galaxy Density Peak
In Fig. 8 we explore the stellar mass and SFR completeness limits required to recover the baseline highest galaxy density peak. The left column shows the 3D separation between the baseline and stellar mass-limited highest galaxy density peaks, with the top, middle, and bottom rows showing the data binned by cluster halo mass at $z=0$. Separations are plotted as a function of redshift, with three stellar mass thresholds, ranging from $10^{9.0}$ to $10^{10.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. While the scatter is significant, we find that the median separations do not generally correlate with the cluster halo mass at $z=0$, with these separations being comparable across $M_{\rm{200}}^{z=0}$ bins. However, there is a correlation with the the minimum stellar mass threshold, with more restrictive thresholds yielding larger median separations. This result highlights the sensitivity of the inferred galaxy density field to the underlying stellar mass completeness limit.
@@
In the right column of Fig. 8, we show the 3D separation between the baseline and SFR-limited highest galaxy density peaks, with the SFR-limited sample restricted to ${M}_{\star}>10^{9.0}\leavevmode\nobreak\ {\rm M}_{\odot}$ and ranging from $0.1$ to $5.0\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$. Once again, we find that the correlation between the cluster halo mass at $z=0$ and the magnitude of the median separations is relatively weak. Additionally, while the scatter is large, the median separations are generally less than $0.5$ pMpc at $z>2$ and do not strongly correlate with the star formation rate thresholds explored in this study.
@@
In Fig. 9, we plot the baseline highest density peak recovery fraction as a function of highest density peak-centric separation. The left column shows the cumulative recovery fraction, binned by redshift from $z=1$ to $z=5$, for the separation between the baseline highest galaxy density peak and ${M}_{\star}$-limited highest galaxy density peaks for stellar mass thresholds ranging from $10^{9.0}$ to $10^{10.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. For ${M}_{\star}>10^{9.0}\leavevmode\nobreak\ {M}_{\star}$, approximately $65-70\%$ of the sample recovers the baseline density peak for separations less than $1$ pMpc; however, this drops to $40-45\%$ for peaks traced by protocluster populations with ${M}_{\star}>10^{10.5}\leavevmode\nobreak\ {\rm M}_{\odot}$.
@@
The right column of Fig. 9 shows the cumulative recovery fraction for the separation between the baseline highest galaxy density peak and the highest galaxy density peaks traced by SFR-limited subpopulations with SFR thresholds ranging from $0.1$ to $5.0\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$. For all of the SFR thresholds explored, the recovery rate is consistently around $65-70\%$ at $z>2$ for scales less than 1.0 pMpc. Given that this sample is constrained to ${M}_{\star}>10^{9.0}\leavevmode\nobreak\ {\rm M}_{\odot}$, this suggests that stellar mass, rather than SFR, is the primary driver of the discrepancy between the location of the SFR-limited and baseline highest galaxy density peaks.
@@
Figure 10: Ratio of the average SFR (top row) and stellar mass (bottom row) within $R_{10}$ centered on the baseline (${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$, red) and ${M}_{\star}$-limited (${M}_{\star}>10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$, blue) highest galaxy density peak, compared to the average SFR and stellar mass in the surrounding volume outside of $R_{10}$. The columns show these ratios binned by cluster mass at $z=0$. A box-and-whisker plot visualizes the distribution, with vertical boxes representing the interquartile range (25th to 75th percentiles), horizontal bars and circles indicating the median and mean ratios, and whiskers extending to 1.5 times the interquartile range, covering approximately $95\%$ of the distribution. The principal difference between the ratio distributions measured relative to the baseline and ${M}_{\star}$-limited peaks is that the former is typically more enhanced than the latter. Nevertheless, both scenarios are consistent with stellar mass growth and star formation activity being enhanced in the inferred densest region relative to the remainder of the protocluster.
@@
To examine the impact of potentially missing the highest galaxy density peak in observations, we explore the properties of galaxies (e.g., SFRs or stellar masses) inside and outside the densest region in our simulated protocluster population. We compare these properties by calculating the ratio of the average SFR and stellar mass inside and outside $R_{10}$, measured relative to the location of the highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$). Ratios close to unity imply that the galaxies in the highest density peak have properties on average similar to those of galaxies in the rest of the protocluster. Ratios below (above) unity indicate that the galaxies within the highest density peak have properties that are depressed (elevated) relative to the remainder of the protocluster.
@@
Fig. 10 shows the ratio of the average SFR (top row) and stellar mass (bottom row) within $R_{10}$ centered on either the baseline (${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$, red) or ${M}_{\star}$-limited (${M}_{\star}>10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$, blue) highest galaxy density peak to the average SFR and stellar mass beyond this region, as a function of redshift. As described in §2.4, $R_{10}$ is measured relative to the center of mass of the baseline protocluster protocluster. In this way, the baseline population can be compared to a stellar mass limited population that is consistent with observational completeness limits. The data is binned by cluster halo mass at $z=0$ and presented as a box-and-whisker plot, with the vertical boxes representing the interquartile range (25th to 75th percentiles), and the horizontal bars and circles showing the median and mean of the distribution. The whiskers extend to 1.5 times the interquartile range, encompassing approximately $95\%$ of the distribution.
We find that the SFR ratios exhibit a mild dependence on the cluster halo mass. At early times, the progenitors of the most massive clusters tend to have higher median SFR ratios than those of less massive systems. However, by $z<3$, this trend reverses, with the most massive progenitors showing lower median SFR ratios. This suggests that, relative to the progenitors of the least massive clusters, star formation is suppressed earlier in the densest regions of the progenitors of the most massive clusters. This is possibly a sign of cluster scale “downsizing” (Cowie et al., 1996; Bower et al., 2006; Fontanot et al., 2009; Oser et al., 2010; Rennehan et al., 2020), in which the galaxies in the densest regions of the progenitors of the most massive clusters preferentially have their star formation suppressed at earlier times.
While the full distributions of average SFR and stellar mass ratios at $z>2$ extend below unity, both the mean and median values are consistently above unity, regardless of the cluster halo mass at $z=0$. These results are unchanged even when using the median, rather than the mean, SFR and stellar mass within and beyond the highest density peak. These ratios suggest that the densest regions of galaxy protoclusters are generally sites of enhanced star formation and rapid galaxy growth, consistent with many observational studies (e.g., Steidel et al., 2005; Dannerbauer et al., 2014; Hayashi et al., 2016; Miller et al., 2018; Shimakawa et al., 2018; Koyama et al., 2021; Polletta et al., 2021). Additionally, while less common, these ratios also support scenarios in which the densest region of galaxy protoclusters exhibit depressed star formation activity, which is consistent with observational findings from Mei et al. (2023) showing that the passive-density relation was already established in some protoclusters at $z\sim 2$.
Considering that our baseline simulated protocluster population extends to galaxies much less massive than those detectable in current surveys suggests that the enhancement of stellar mass growth and star formation in densest regions of protoclusters is insensitive to the stellar mass completeness limits explored in this study. However, as shown by comparing the baseline (red) and ${M}_{\star}$-limited (blue) ratio distributions in Fig. 10, we see that while the median and mean ratios are typically above unity, the baseline ratio distribution is systematically higher. This indicates that the inclusion of low-mass galaxies in observations would further boost the commonly observed enhancement of star formation activity and stellar mass growth in the densest regions of protoclusters.
Figure 11: Projected density distribution for the progenitor of the most massive $z=0$ galaxy cluster as a function of redshift. The top row shows the galaxy density map, while the middle and bottom rows display the stellar mass and dark matter weighted density maps from TNG-Cluster, respectively. The location of the highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$) is marked by a black cross, while the highest stellar mass ($\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$) and dark matter ($\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$) peaks are indicated by black plus and hexagon symbols, respectively. For this specific protocluster, we find that the galaxy density peaks generally coincide with the highest concentrations of stellar mass and dark matter, with the only exception occurring at $z=3$ for the stellar mass density peak.
@@
Figure 12: Cumulative recovery fraction of the baseline stellar mass and dark matter highest density peaks, binned by redshift and measured relative to the highest galaxy density peaks traced by various ${M}_{\star}$- and SFR-limited subpopulations. The first two rows show separations from the baseline stellar mass (first row) and dark matter (second row) highest density peaks to the highest galaxy density peak traced by ${M}_{\star}$-limited subpopulations. The third and fourth rows show the same for SFR-limited subpopulations. Columns represent increasingly restrictive stellar mass ($10^{8.5}$ to $10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$) and SFR (0.10.10.10.$10.0\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$) thresholds. For the least restrictive SFR and stellar mass thresholds, we find that for $z>2$ the highest stellar mass (dark matter) density peaks are recovered in $\lesssim 60-65\%$ ($\lesssim 75-80\%$) of cases with accuracies within 1.0 pMpc ($2.1-2.6\arcmin$).
@@
Thus far we have considered only the region of the highest galaxy density peak, as galaxy number densities are the primary method in which the density field of observed protoclusters are characterized. However, the highest matter density peak could be examined by weighting the galaxy density field according to the stellar and halo masses of the protocluster galaxies. Comparing the matter density peaks with the galaxy density peaks allows us to ascertain if the highest concentrations of galaxies is consistent with the highest concentration of luminous and dark matter. To achieve a balanced weighting scheme that prevents either high-mass or low-mass galaxies from dominating the density field, we define the weights applied to the 3D histogram as $w=M^{\alpha}+C$, where $M$ represents either the stellar mass or dark matter component of each galaxy. We set $\alpha=0.5$ to ensure a balanced contribution from low-mass and high-mass galaxies, while the constant $C=0.5$ establishes a baseline weight [^2] In practice, $\alpha$ only affects the map in the range between 0 and 1, with $\alpha\leq 0$ returning the original galaxy density map, and $\alpha\geq 1$ excluding contributions from low-mass galaxies, making the map solely reflect the location of the most massive galaxy..
@@
In Fig. 11 we present the projected galaxy density (top row), stellar mass density (middle row), and halo mass density (bottom row) for the progenitor of the most massive galaxy cluster at $z=0$. The columns display the projected density maps at five distinct redshifts, ranging from $z=1$ to $z=5$. We overlay the location of the highest density peak for each density field, marking the highest galaxy density peak with a black cross, the highest stellar mass peak with a black plus sign, and the highest dark matter mass density peak with a black hexagon. For this specific test case we find that the location of the highest galaxy density peak almost always coincides with the location of the highest stellar mass and dark matter density peaks, with the only exception occurring at $z=3$ for the stellar mass density peak.
@@
Figure 13: Box-and-whisker plot of the ratio of the average SFR (top row) and stellar mass (bottom row) within $R_{10}$ centered on the highest density peak for dark matter (blue boxes, $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$), stellar mass (orange boxes, $\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$), and galaxies (yellow boxes, $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$), compared to the average SFR and stellar mass in the surrounding volume outside $R_{10}$. The columns display these ratios in bins of cluster mass at $z=0$. Regardless of whether the measurements are made relative to galaxy or matter density peaks, the SFR and stellar mass ratios remain above unity at $z>2$. The primary difference is that ratios measured relative to matter density peaks are systematically higher.
@@
The first two rows of Fig. 12 show the cumulative matter density recovery fractions measured from the baseline stellar mass (first row) and dark matter (second row) highest density peaks to the highest galaxy density peak traced by four ${M}_{\star}$-limited subpopulations. Moving from left to right, the stellar mass thresholds increase from $10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ (i.e., the baseline stellar mass established in this study) to $10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$. The highest stellar mass density peaks are recovered in $\lesssim 60-65\%$ of cases within an accuracy of 1.0 pMpc ($\sim 2.1-2.6\arcmin$) at $2<z\leq 5$, with only a weak dependence on the minimum stellar mass threshold. Alternatively, the dark matter density peaks are recovered in $\lesssim 75-80\%$ of the cases within an accuracy of 1.0 pMpc at $2<z\leq 5$ for a minimum stellar mass threshold of ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$, but within the same accuracy and redshift redshift range the recovery fraction drops to $\lesssim 50-60\%$ when the minimum stellar mass threshold is raised to ${M}_{\star}>10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$.
@@
The last two rows of Fig. 12 show the cumulative matter density recovery fractions measured from the baseline stellar mass (third row) and dark matter (fourth row) highest density peaks to the highest galaxy density peak traced by four SFR-limited subpopulations. Moving from left to right, the SFR thresholds increase from $0.1$ to $10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$. We find that matter density peaks are recovered in $\lesssim 60-70\%$ of cases within an accuracy of $1.0$ pMpc ($\sim 2.1-2.6\arcmin$) for $2<z\leq 5$.
The specific recovery fractions quoted above depend on the weighting scheme used to construct the matter density maps. We adopt $\alpha=0.5$ as it offers a balanced compromise between $\alpha\leq 0$, which reproduces the baseline galaxy density field, and $\alpha\geq 1$, which places disproportionate weight on the location of the most massive galaxy. Small deviations from this parameter choice have only a marginal effect on the agreement between the galaxy and matter density peaks. Despite this caveat, the results suggest that the highest-density peaks traced by galaxies generally coincide with those of the underlying matter distribution.